Solve this system for (x, y) : 2^x.4^y = 16 and 4^3x-2.(1/2)^y = 128
Answer by vksarvepalli(154)

2^x.4^y = 16
so 2^x.2^2y = 2^4
=> 2^(x+2y) = 2^4
=> x+2y=4 -------------- 1
4^3x-2.(1/2)^y = 128
=> 2^(6x-4).2^(-y) = 2^7
=> 6x-4-y = 7
=> 6x-y=11 -------------- 2
solving equations 1 and 2 (multiply eq. 2 by 2 and add eq.1 )
x=2
y=1